OpenStudy (tumblewolf):

Quick dilution questions!! Will fan and medal!! 1) It takes how many mL of 10mg/mL Sucrose solution to get 10mL of 1mg/mL. For this one I got the answer being 1 mL because (v1)(c1)=(v2)(c2). The answer seems too easy and I feel like it is trying to trip me up. 2) How many mg of NaCl is required to make 0.1 L of 5mg/mL salt solution. I think the wording is throwing me off. I'm tired and my brain isn't switching the words around like it usually does. Thank you ahead of time!!

OpenStudy (photon336):

so, you're starting with 10 mg/mL and you want to dilute it down to 1~mg/mL

OpenStudy (photon336):

\[v_{1}c_{1} = v_{2}c_{2}\] so you would solve for v1 the initial volume you need. you know your final volume. \[\frac{ v_{2}c_{2} }{ c_{1} } = v_{1}\] now plug in \[\frac{ 10~\cancel\mL*(\frac{ \cancel\mg }{ \cancel\mL }) }{ 10~(\frac{ \cancel\mg }{ mL }) } = \frac{ 10 }{ 10*(\frac{ 1 }{ mL }) } = 1~mL ~v_{1}\]

OpenStudy (photon336):

what's \[0.1~\cancel\L ~ *\frac{ (1,000~mL) }{ \cancel\L } = 100~mL\] \[100~\cancel\mL*(\frac{ 5~mg }{ \cancel\mL}) = 500~mg \]